# MixGRPO: Unlocking Flow-based GRPO Efficiency with Mixed ODE-SDE

### v1

@@ Appendix C MixGRPO-Flash Algorithm
9:   for generate $i$-th image from $i=1$ todo
@@
19:     Calculate multi-reward advantage: $A_{i}\leftarrow\sum_{k=1}^{K}\frac{R(\mathbf{x}^{i}_{\tilde{T}},\mathbf{c})^{i}_{k}-\mu_{k}}{\sigma_{k}}$
20:   end for
23:   end for
24:  end for
25:  if use MixGRPO-Flash* then $\triangleright$ move sliding window
26:   $l\leftarrow 0$
27:  else
28:   if $m\bmod\tau~\text{is}~0$ then
29:     $l\leftarrow\min(l+s,~T-w)$
30:   end if
31:  end if
32:end for
$S=\frac{T}{w+(T-w)\tilde{r}}$ (18)
$S=\frac{T}{\mathbb{E}_{l}\left(w+l+\lceil(T-w-l)\tilde{r}\rceil\right)}<\frac{T}{w+(T-w)\tilde{r}}$ (19)
Appendix D More visualized results
@@
Figure 7: Comparison of the visualization results of FLUX, DanceGRPO, and MixGRPO under HPS-v2.1 and CLIP Score as the reward models.

### v2

@@ Appendix D MixGRPO Algorithm
9:   for generate $i$-th image from $i=1$ to N do
10:     for sampling timestep $t=0$ to T1T-1 do $\triangleright$ $\pi_{\theta_{\text{old}}}$ mixed sampling loop
11:      if $t\in W(l)$ then
12:        Use SDE Sampling to get $\mathbf{x}^{i}_{t+1}$
13:      else
14:        Use ODE Sampling to get $\mathbf{x}^{i}_{t+1}$
15:      end if
16:     end for
17:   end for
18:   for $i$-th image from $i=1$ to N do
19:     Calculate multi-reward advantage: $A_{i}\leftarrow\sum_{k=1}^{K}\frac{R(\mathbf{x}^{i}_{T},\mathbf{c})^{i}_{k}-\mu_{k}}{\sigma_{k}}$
20:   end for
23:   end for
24:  end for
25:  if $m\bmod\tau~\text{is}~0$ then $\triangleright$ move sliding window
26:   $l\leftarrow\min(l+s,~T-w)$
27:  end if
28:end for
Appendix E MixGRPO-Flash Algorithm
MixGRPO-Flash Algorithm 2 accelerates the ODE sampling that does not contribute to the calculation of the policy ratio after the sliding window by using DPM-Solver++ in the Eq. (21). We introduce a compression rate $\tilde{r}$ such that the ODE sampling after the window only requires $(T-l-w)\tilde{r}$ time steps. And the total time-steps is $\tilde{T}=l+w+(T-l-w)\tilde{r}$ The final algorithm is as follows:
$S=\frac{T}{w+(T-w)\tilde{r}}$. (22)
$S=\frac{T}{\mathbb{E}_{l}\left(w+l+\lceil(T-w-l)\tilde{r}\rceil\right)}<\frac{T}{w+(T-w)\tilde{r}}$. (23)
Algorithm 2 MixGRPO-Flash Training Process
1:initial policy model $\pi_{\theta}$; reward models $\{R_{k}\}_{k=1}^{K}$; prompt dataset $\mathcal{C}$; total sampling steps $\tilde{T}$; number of samples per prompt NN; ODE compression rate $\tilde{r}$
3:Init left boundary of $W(l)$: $l\leftarrow 0$
4:for training iteration $m=1$ to $M$ do
5:  Sample batch prompts $\mathcal{C}_{b}\sim\mathcal{C}$
7:  for each prompt $\mathbf{c}\in\mathcal{C}_{b}$ do
8:   Init the same noise $\mathbf{x_{0}}\sim\mathcal{N}(0,\mathbf{I})$
9:   for generate $i$-th image from $i=1$ to N do
@@ Appendix E MixGRPO-Flash Algorithm
19:   end for
20:   for $i$-th image from $i=1$ to N do
21:     Calculate multi-reward advantage: $A_{i}\leftarrow\sum_{k=1}^{K}\frac{R(\mathbf{x}^{i}_{\tilde{T}},\mathbf{c})^{i}_{k}-\mu_{k}}{\sigma_{k}}$
22:   end for
23:   for optimization timestep $t\in W(l)$ do $\triangleright$ optimize policy model $\pi_{\theta}$
24:     Update policy model via gradient ascent: $\theta\leftarrow\theta+\eta\nabla_{\theta}\mathcal{J}$
25:   end for
26:  end for
27:  if use MixGRPO-Flash* then $\triangleright$ move sliding window
28:   $l\leftarrow 0$
29:  else
30:   if $m\bmod\tau~\text{is}~0$ then
31:     $l\leftarrow\min(l+s,~T-w)$
32:   end if
33:  end if
34:end for
Appendix F More visualized results
@@
Figure 7: Comparison of the visualization results of FLUX, DanceGRPO, and MixGRPO under HPS-v2.1 and CLIP Score as multi-reward models.
Figure 8: Comparison of the visualization results of SD3.5-M, offline DPO, online DPO, Flow-GRPO and MixGRPO under HPS-v2.1, Pick Score and ImageReward as multi-reward models.